find the value of k for which the following equation has two equal roots.

(k+2)x^2+4k=(4k+2)x


1
Expert's answer
2022-01-12T04:22:12-0500

Compute the discriminant of the quadratic "a x^{2}+b x+c=0."

It is:

"\Delta=b^{2}-4 a c=(4 k+2)^{2}-4 \cdot 4 k(k+2)\\=16 k^{2}+16 k+4-16 k^{2}-32 k=-16 k+4"

A quadratic has two equal roots if and only if "\Delta=0\\"

"-16k+4=0\Rightarrow k=\frac{4}{16}=\frac{1}{4}"

Hence, the value of k is "\frac{1}{4}" .